# PDNNet: PDN-Aware GNN-CNN Heterogeneous Network for Dynamic IR Drop Prediction

### v1

@@ Abstract
IR drop on the power delivery network (PDN) is closely related to PDN’s configuration and cell current consumption. As the integrated circuit (IC) design is growing larger, dynamic IR drop simulation becomes computationally unaffordable and machine learning based IR drop prediction has been explored as a promising solution. Although CNN-based methods have been adapted to IR drop prediction task in several works, the shortcomings of overlooking PDN configuration is non-negligible. In this paper, we consider not only how to properly represent cell-PDN relation, but also how to model IR drop following its physical nature in the feature aggregation procedure. Thus, we propose a novel graph structure, PDNGraph, to unify the representations of the PDN structure and the fine-grained cell-PDN relation. We further propose a dual-branch heterogeneous network, PDNNet, incorporating two parallel GNN-CNN branches to favorably capture the above features during the learning process. Several key designs are presented to make the dynamic IR drop prediction highly effective and interpretable. We are the first work to apply graph structure to deep-learning based dynamic IR drop prediction method. Experiments show that PDNNet outperforms the state-of-the-art CNN-based methods by up to $39.3\%$ reduction in prediction error and achieves 545×545$\times$545 × speedup compared to the commercial tool, which demonstrates the superiority of our method.
@@ I Introduction
In literature, convolutional neural networks (CNNs) in machine learning have dominated this task for a long time. Current CNN-based methods regard dynamic IR drop prediction as an image generation process [12, 13, 14, 15, 16]. Mozaffari et al. [17] develop an on-chip power supply noise evaluation method, leveraging the CNN model to handle global and local features in prediction. Xie et al. [18] propose a max-CNN structure, called PowerNet, that encodes input power features to capture the peak voltage drop along the temporal axis. Chhabria et al. [19] propose MAVIREC, which formulates the prediction task as an image-to-image translation problem, utilizing a UNet architecture to percept dynamic current variation.
Although these methods achieve good performance, the prominent drawbacks can not be neglected. As Figure 1 shows, current load patterns and PDN architectures are two indispensable parts that jointly impact the final IR drop distribution. Conventional works focus on extracting the representation from current load patterns, while almost none of them devote attention to explore the PDN structures. The reason mainly lies in two-fold. Firstly, image-like input features cannot fully represent PDN structure and current flow. Secondly, as pure CNN-based models lack topology relation aggregation approaches, they are neither capable of perceiving fine-grained cell-PDN connection relation, nor modeling the current flowing behavior on PDN.
@@
  • We conduct extensive experiments on a large-scale dataset. PDNNet outperforms the state-of-the-art CNN-based methods by up to 39.3% reduction in prediction error and achieves 545$\times$ speedup compared to the commercial tool, which demonstrates the superiority of our proposed method.
@@ II-A IR Drop Analysis
IR drop analysis simulates the voltage drop under the combined effect of parasitic and current flow through the PDN [22]. Generally, PDN is casted as a grid topology network with parameters, that mainly include impedance and conductance.[23, 24, 25, 26, 27, 28, 29, 30, 31] Voltage sources (i.e., power supply rails) and current loads (i.e., cells) respond to simulate the current flowing on the network. The illustration of PDN grid depicts in Figure 2.
@@ III-A1 Graph Construction
We divide the whole layout into uniform grids. The Figure 4 (a) illustrates our PDNGraph outline, where we define the part within a grid as the node of PDNGraph, while the edge connects adjacent nodes in one-hop is restricted only exists in the horizontal and vertical directions. The inter-node distance, corresponding to the resistance distribution, is also assumed uniform across the whole design [18]. In PDNGraph, we opt not to explicitly calculate resistance and capacitance for each edge, for fully believing electrical law can be “learned” by an artificial neural network (e.g., our PDNNet). Hereafter, PDNGraph manifests a consistent manner with the grid network in IR drop analysis.
@@ III-A2 Current Load Attribute
We compose and assign these features to form node features $x$, $x\in\mathbb{R}{{}^{N\times C}}$, where $N$ denotes the node number in PDNGraph and $C$ represents the input feature dimension.
@@ III-A3 Current Direction Attribute
In the general PDN configuration, the high power supply VDD rails (HPR) transmit current to the low redistribution VDD layer (LRL). Despite taking complete PDN into account, we opt for the most relevant part of IR drop distribution to avoid redundancy. Firstly, without considering the macros, LRL (i.e., M0 or M1) is the layer that directly connects to the cells where most of the IR drop occurs. Secondly, the location where crisscrossed HPR (i.e., M7 in  [33]) directly connects to LRL through vias has the least IR drop violation. To this end, the power supply process can be illustrated that the current flow spread from HPR ends to LRL, along the LRL strip direction towards each node. The voltage drop at this point, as visually explained in Figure 4 (b)(c), also increases gradually when LRL moves away from the HPR. Therefore, we designated the edges’ direction starting from the closest HPR towards the away direction. Since there is no explicit current direction between vertical nodes, we consider them as undirectional edges. Further, we neglect the influence of current on the voltage drop across HPR with respect to their low resistance nature. Notice that our PDNGraph can be adaptive to the uniform or non-uniform PDN with high flexibility, as shown in Figure 4 (a).
@@ III-B1 GNN Branch
Voltage-Drop Acquiring Layer. To utilize the current-related features to learn the voltage drop, we stack $l$ voltage-drop acquiring blocks to form this layer. To better acquire one nodei’s voltage drop representation, we sum all the neighbors’ voltage drop features,
$m_{i}\ ={\displaystyle\sum_{k\in n(i)}\left(v_{k}^{(l-1)}\right)},$ (3)
$n$ is nodei’s neighbors and $m_{i}$ is the neighbors’ feature summation. Therefore, we update the feature of this node by concatenating all neighbor effects,
$v_{i}^{(l)}=f_{\phi}(v_{i}^{(l-1)}\ \|\ m_{i}),$ (4)
@@
Neighbor Influence Layer. Since the node features have already captured the characteristics of voltage drop through the above layers, the objective of this layer is to learn and refine the influence of neighbors on each node. We therefore update the PDNGraph to an undirected graph while still preserving its topological structure. Subsequently, learnable function blocks $f_{\theta}$ are employed to update the node representation using both the original node features and the ones from its neighbors,
$v_{i}^{(l)}=f_{\theta}(v_{i}^{(l-1)}\ \|\ v_{k}^{(l-1)})\ \ k\in n(i),$ (5)
@@ III-B2 CNN Branch
CNN branch in PDNNet is performed as a parallel temporal branch of GNN’s. Although GNN-branch can obtain considerable performance, CNN branch is indispensable for perceiving current alteration along the time axis in dynamic IR drop prediction. Details can be founded in the experiment Section IV. To achieve temporal perception, we directly adopt a classical encoder-decoder structure with 2 shortcut connections. The architecture is present in the bottom of Figure 5. More advanced architectures can be applied here, but this is not the focus of this work. Details of the CNN branch are described as follows. The intake features share the same as those of the GNN branch. We map the graph node features onto a 2D space and insert a dummy dimension instead of the original channel. The transformed feature shape is $x\in\mathbb{R}{{}^{H\times W\times T\times C}}$, where $T$ denotes the number of input features and $C$ means feature channels through the representation process. The input feature is first fed into a 3D encoder network that goes through $l$ encoder layers, each layer comprise of two 3D convolutional and down-sampling blocks. After feature encoding, the output embedding shape converts to $x^{cnn}\in\mathbb{R}{{}^{\frac{H}{2l}\times\frac{W}{2l}\times\frac{T}{l}\times C}}$, where $l$ opt as 3 in our model. At this point, we shrink the temporal channels since they have been sufficiently extracted.
$x^{cnn}\in\mathbb{R}{{}^{\frac{H}{2l}\times\frac{W}{2l}\times\frac{T}{l}\times C}}\ \longrightarrow\ x^{cnn}\in\mathbb{R}{{}^{\frac{H}{2l}\times\frac{W}{2l}\times C}}$ (7)
Subsequently, comprising $l$ layers of deconvolution and up-sampling blocks, the 2D decoder network maintains mirror symmetry in spatial shape with the 3D encoder counterpart at each level. To further assist the recovery of low-level information from the original feature, we fuse the 4D features, as shown in Figure 5, from the left side into each layer’s 2D block. Finally, the CNN network outputs IR drop temporal representation $y^{cnn}\in\mathbb{R}{{}^{H\times W\times 1}}$.
TABLE II: Evaluation results on CircuitNet dataset. Our PDNNet achieves consistent performance gain on five evaluation metrics in both large and small datasets. Red denotes percentage of improvement. $\downarrow$ means “lower is better”, $\uparrow$ means “higher is better”.
Method CircuitNet [33]
[…]
PowerNet [18] 0.107 0.59 0.74 0.72 0.62
PDNNet (Ours) 0.031(-71.0%) 0.90(+52.5%) 0.95(+28%) 0.75(+4.2%) 0.63(+1.6%)
MAVIREC [19] 0.038 0.47 0.89 0.74 0.62
PDNNet (Ours) 0.031(-18.4%) 0.90(+91.5%) 0.95(+6.7%) 0.75(+1.4%) 0.63(+1.6%)
@@ III-C Fusion Module
In the last part, we fuse upstream embedding together to get a comprehensive representation. We first align the feature dimensions by flattening the temporal representation $y^{cnn}\in\mathbb{R}{{}^{H\times W\times 1}}$ into $y^{cnn}\in\mathbb{R}{{}^{HW}}$. Then, a concatenation function is imposed to merge these features and pass them through an $MLP$ layer as a regressor to generate the final IR drop prediction.
$z^{prediction}\ =\ MLP(y^{gnn}\ \ \|\ \ y^{cnn}),$ (8)
@@ III-D Discussion
We notice that some literature attempts to involve PDN features. The previous work [34] targets static IR drop prediction that regards the PDN as a 2D density map, where each pixel denotes the total number of PDN strips covering this location. However, our PDNNet manifests an obvious difference with this method. First, our method targets more difficult dynamic IR drop prediction. Second, we construct a graph structure based on the intuition of IR drop analysis to finely represent the cell-PDN physical relation, compared to the image-like PDN density map. Third and finally, we design a dual-branch GNN-CNN heterogeneous network to favorably aggregate PDN attributes and temporal variation for accurate dynamic IR drop prediction, compared to vanilla image encoder-decoder network (like Figure 3 (b)).

### v2

@@ Abstract
IR drop on the power delivery network (PDN) is closely related to PDN’s configuration and cell current consumption. As the integrated circuit (IC) design is growing larger, dynamic IR drop simulation becomes computationally unaffordable and machine learning based IR drop prediction has been explored as a promising solution. Although CNN-based methods have been adapted to IR drop prediction task in several works, the shortcomings of overlooking PDN configuration is non-negligible. In this paper, we consider not only how to properly represent cell-PDN relation, but also how to model IR drop following its physical nature in the feature aggregation procedure. Thus, we propose a novel graph structure, PDNGraph, to unify the representations of the PDN structure and the fine-grained cell-PDN relation. We further propose a dual-branch heterogeneous network, PDNNet, incorporating two parallel GNN-CNN branches to favorably capture the above features during the learning process. Several key designs are presented to make the dynamic IR drop prediction highly effective and interpretable. We are the first work to apply graph structure to deep-learning based dynamic IR drop prediction method. Experiments show that PDNNet outperforms the state-of-the-art CNN-based methods and achieves 545×545$\times$545 × speedup compared to the commercial tool, which demonstrates the superiority of our method.
@@ I Introduction
In literature, convolutional neural networks (CNNs) in machine learning have dominated this task for a long time. Current CNN-based methods regard dynamic IR drop prediction as an image generation process [12, 13, 14, 15, 16, 17]. Mozaffari et al. [18] develop an on-chip power supply noise evaluation method, leveraging the CNN model to handle global and local features in prediction. Xie et al. [19] propose a max-CNN structure, called PowerNet, that encodes input power features to capture the peak voltage drop along the temporal axis. Chhabria et al. [20] propose MAVIREC, which formulates the prediction task as an image-to-image translation problem, utilizing a UNet architecture to percept dynamic current variation.
Although these methods achieve good performance, the prominent drawbacks can not be neglected. As Figure 1 shows, current load patterns and PDN architectures are two indispensable parts that jointly impact the final IR drop distribution. Study [21] uses an image-like PDN density map to represent cross-grained PDN structure for static IR drop prediction. However, in the dynamic IR drop prediction field, conventional works focus on extracting the representation from current load patterns, while almost none of them devote attention to exploring the PDN structures. The reason mainly lies in two-fold. Firstly, image-like input features cannot fully represent PDN structure and dynamic current flow, especially powerless when facing the fairly regular PDN mesh (i.e. PDN density map has the same value on each pixel in this case). Secondly, as pure CNN-based models lack topology relation aggregation approaches, they are neither capable of perceiving fine-grained cell-PDN connection relation, nor modeling the current flowing behavior on PDN.
@@
  • We conduct extensive experiments on a large-scale dataset. PDNNet outperforms the state-of-the-art CNN-based methods by up to 24.3% reduction in prediction error and achieves 545$\times$ speedup compared to the commercial tool, which demonstrates the superiority of our proposed method.
@@ II-A IR Drop Analysis
IR drop analysis simulates the voltage drop under the combined effect of parasitic and current flow through the PDN [24]. Generally, PDN is casted as a grid topology network with parameters, that mainly include impedance and conductance.[25, 26, 27, 28, 29, 30, 31, 32] Voltage sources (i.e., power supply rails) and current loads (i.e., cells) respond to simulate the current flowing on the network. The illustration of PDN grid depicts in Figure 2.
@@ III Method
Figure 5: The figure shows how the directions of edges between nodes in PDNGraph are determined. We categorize the relationship between the PDN location and the current node into four cases.
@@ III-A PDNGraph
Algorithm 1 PDNGraph Construction Algorithm
0:  Layout dimensions $W$ and $H$, grid $g$ with fixed size ($d^{x}$, $d^{y}$), a set of vertical PDN with $x$ axis position $\{PDN^{x}\}$.
0:  Our PDNGraph.
1:  Divide the layout using grid $g$, $N_{w}=\frac{W}{d^{x}},N_{h}=\frac{H}{d^{y}}$.
2:  for $i=1$ to $N_{h}$ do
3:     for $j=1$ to $N_{w}$ do
4:        Set the initial value $s=\infty$.
5:        for each $PDN^{x}_{k}$ in $\{PDN^{x}\}$ do
6:           Compute the coordinate subtraction $s$ between $PDN_{k}$ and current node $v_{i,j}$, $s\_tmp=PDN^{x}_{k}-v_{i,j}^{x}$.
7:           if $s\_tmp<s$ then
8:              $s=s\_tmp$
9:           end if
10:        end for
11:        if $j<N_{w}$ then
12:           if $s\leq 0$ then
13:              Set the edge direction from node $v_{i,j}$ to $v_{i,j+1}$ (the closest $PDN$ is on the left of the current node $v_{i,j}$), edge is $v_{i,j}\longrightarrow v_{i,j+1}$.
14:           else if $0<s\leq 1$ then
15:              Set the bidirectional edge between node $v_{i,j}$ and $v_{i,j+1}$ (the closest $PDN$ is between node $v_{i,j}$ and $v_{i,j+1}$), edge is $v_{i,j}\longleftrightarrow v_{i,j+1}$.
16:           else if $1<s$ then
17:              Set the edge direction from node $v_{i,j+1}$ to $v_{i,j}$ (the closest $PDN$ is on the right of the current node $v_{i,j}$ and neighbor node $v_{i,j+1}$), edge is $v_{i,j+1}\longrightarrow v_{i,j}$.
18:           end if
19:        end if
20:        if $i<N_{h}$ then
21:           Set the bidirectional edge between node $v_{i,j}$ and $v_{i+1,j}$, edge is $v_{i,j}\longleftrightarrow v_{i+1,j}$.
22:        end if
23:     end for
24:  end for
@@ III-A1 Graph Construction
The Figure 4 (a) illustrates our PDNGraph construction outline. We regard the circuit layout as a checkerboard, defining each grid $g$ that covers multiple cells as a “node” $v$ of our PDNGraph. The edges connect adjacent nodes $N_{v}$ in one-hop and are restricted only in the horizontal and vertical directions, depicted in Figure 4 (b)(c). The edge distances between nodes implicitly encapsulate the resistance and capacitance characteristics of the power supply metal. To ensure uniformity in these attributes, we employ a “uniform grid” strategy same as [19] for constructing the PDNGraph, which guarantees consistent edge properties across the graph. Assuming the layout size is $W\times H$, we utilize fixed grids size $d_{x}\times d_{y}$ to segment the layout to $N_{w}\times N_{h}$, where $N_{w}=\frac{W}{d_{x}}$ and $N_{h}=\frac{H}{d_{y}}$. In PDNGraph, we opt not to explicitly calculate resistance and capacitance for each edge, for fully believing electrical law can be “learned” by an artificial neural network (e.g., our PDNNet). Hereafter, PDNGraph manifests a consistent manner with the grid network in IR drop analysis. Note that the grid division is not dependent on the PDN structure and would only affect the edge direction in the PDNGraph that is elaborated in the current direction attribute part.
Figure 6: (a) message aggregation and passing on PDNGraph. (b) PDNGraph construction of regular PDN grid. (c) PDNGraph construction of irregular PDN grid. Generally, standard cells that are further away from the PDN experience higher IR drop.
@@ III-A2 Current Load Attribute
We sum each cell power feature within a gird to form the overall node features $x$, $x\in\mathbb{R}{{}^{N\times C}}$, where $N$ denotes the node number in PDNGraph, $N=N_{w}\times N_{h}$ and $C$ represents the input feature dimension.
@@ III-A3 Current Direction Attribute
In the general PDN configuration, the high power supply VDD rails (HPR) transmit current to the low redistribution VDD layer (LRL). Despite taking complete PDN into account, we opt for the most relevant part of IR drop distribution to avoid redundancy. Firstly, without considering the macros, LRL (i.e., M0 or M1) is the layer that directly connects to the cells where most of the IR drop occurs. Secondly, the location where crisscrossed HPR directly connects to LRL through vias has the least IR drop violation. To this end, the power supply process can be illustrated that the current flow spread from HPR ends to LRL, along the LRL strip direction towards each node. The voltage drop at this point, as visually explained in Figure 4 (b)(c), also increases gradually when LRL moves away from the HPR. Therefore, we designated the edges’ direction oriented away from the nearest high-layer power strip to represent the directional flow of current through the PDN structure.
Specifically, as illustrated in lines 4-10 of Algorithm 1, PDNGraph primarily utilizes the smallest coordinate subtraction $s$ between set of vertical PDN’s position $\{PDN^{x}\}$ in $x$ axis, and current node position $v_{i,j}^{x}$. We categorize the construction of graph edges into three cases based on $s$, as illustrate in Figure 5. Figure 5 (a)(b) corresponds to lines 12-13, Figure 5 (c) corresponds to lines 14-15, Figure 5 (d) corresponds to lines 16-17, and finally, vertical edge is determined as bidirectional corresponds to lines 20-22 in the Algorithm.
That is, the nodes with the same current pattern, may demonstrate distinct IR drop values at different positions among the two high metal rails. Since there is no explicit current direction between vertical nodes, we consider them as bidirectional edges. Further, we neglect the influence of current on the voltage drop across HPR with respect to their low resistance nature. Note that the edge directions just approximate the current directions. It does not matter that much when the real current directions differ from the edge directions we set, since we allow current values to be negative according to what is learnt from the training data.
Additionally, our PDNGraph can be adaptive to regular and irregular PDN with high flexibility, we illustrate the graph construction for different types of PDN structures through figures, shown in Figure 6 (a)(b). Typically, the IR drop value is significantly influenced by the footprint to the closest high-layer PDN strip. Therefore, areas within a wider PDN pitch are likely to exhibit higher IR drop values compared to those within a narrower pitch, given identical current consumption. Consequently, our PDNGraph employs variable edge directions between nodes, which are determined by the location of HPR to reflect the current flow direction through PDN.
@@ III-B1 GNN Branch
Voltage-Drop Acquiring Layer. To utilize the current-related features to learn the voltage drop, we stack $l$ voltage-drop acquiring blocks to form this layer. To better acquire one node $v_{i,j}$’s voltage drop representation, we sum all the neighbors’ voltage drop features,
$m_{v}\ ={\displaystyle\sum_{k\in N_{v}}\left(v_{k}^{(l-1)}\right)},$ (3)
Node $k$ is within the neighborhood $N_{v}$ of node $v$, $v_{k}^{(l-1)}$ represents the feature of node $k$ at layer $l-1$ and $m_{v}$ is the neighbors’ feature summation. Therefore, we update the feature of this node by concatenating all neighbor effects,
$v_{i,j}^{(l)}=f_{\phi}(v_{i,j}^{(l-1)}\ \|\ m_{v}),$ (4)
@@
Neighbor Influence Layer. Since the node features have already captured the characteristics of voltage drop through the above layers, the objective of this layer is to learn and refine the influence of neighbors on each node. We therefore update the PDNGraph to an bidirected graph while still preserving its topological structure. Subsequently, learnable function blocks $f_{\theta}$ are employed to update the node representation using both the original node features and the ones from its neighbors,
$v_{i,j}^{(l)}=f_{\theta}(v_{i,j}^{(l-1)}\ \|\ v_{k}^{(l-1)})\ \ k\in N_{v},$ (5)
@@
TABLE II: The statistics of CircuitNet. Note that training & testing datasets have no intersection in designs.
{tabu}
l—c—c—ccc Design Number of Samples Netlist Statistics
#Cells #Nets Cell Area ($\mu m^{2}$)
Train RISCY-a 2003 44,836 80,287 65,739
RISCY-b 1858 30,207 58,452 69,779
RISCY-FPU-a 1969 61,677 106,429 75,985
RISCY-FPU-b 1248 47,130 84,676 80,030
Test zero-riscy-a 2042 35,017 67,472 58,631
zero-riscy-b 1122 20,350 45,599 62,648
@@ III-B2 CNN Branch
CNN branch in PDNNet is performed as a parallel temporal branch of GNN’s. Although GNN-branch can obtain considerable performance, CNN branch is indispensable for perceiving current alteration along the time axis in dynamic IR drop prediction. Details can be founded in the experiment Section IV. To achieve temporal perception, we directly adopt a classical encoder-decoder structure with 2 shortcut connections. The architecture is present in the bottom of Figure 7. More advanced architectures can be applied here, but this is not the focus of this work. Details of the CNN branch are described as follows. The intake features share the same as those of the GNN branch but align to a fixed size $H_{f}\times W_{f}$. We map the graph node features onto a 2D space and insert a dummy dimension instead of the original channel. The transformed feature shape is $x\in\mathbb{R}{{}^{H_{f}\times W_{f}\times T\times C}}$, where $T$ denotes the number of input features and $C$ means feature channels through the representation process. The input feature is first fed into a 3D encoder network that goes through $l$ encoder layers, each layer comprise of two 3D convolutional and down-sampling blocks. After feature encoding, the output embedding shape converts to $x^{cnn}\in\mathbb{R}{{}^{\frac{H_{f}}{2l}\times\frac{W_{f}}{2l}\times\frac{T}{l}\times C}}$, where $l$ opt as 3 in our model. At this point, we shrink the temporal channels since they have been sufficiently extracted.
$x^{cnn}\in\mathbb{R}{{}^{\frac{H_{f}}{2l}\times\frac{W_{f}}{2l}\times\frac{T}{l}\times C}}\ \longrightarrow\ x^{cnn}\in\mathbb{R}{{}^{\frac{H_{f}}{2l}\times\frac{W_{f}}{2l}\times C}}$ (7)
Subsequently, comprising $l$ layers of deconvolution and up-sampling blocks, the 2D decoder network maintains mirror symmetry in spatial shape with the 3D encoder counterpart at each level. To further assist the recovery of low-level information from the original feature, we fuse the 4D features, as shown in Figure 7, from the left side into each layer’s 2D block. Finally, the CNN network outputs IR drop temporal representation $y^{cnn}\in\mathbb{R}{{}^{H_{f}\times W_{f}\times 1}}$.
@@ III-C Fusion Module
In the last part, we fuse upstream embedding together to get a comprehensive representation. We first interpolate the feature dimensions from $y^{cnn}\in\mathbb{R}{{}^{H_{f}\times W_{f}\times 1}}$ to $y^{cnn}\in\mathbb{R}{{}^{H\times W\times 1}}$. Then we flattening the temporal representation $y^{cnn}\in\mathbb{R}{{}^{H_{f}\times W_{f}\times 1}}$ into $y^{cnn}\in\mathbb{R}{{}^{H_{f}W_{f}}}$. A concatenation function is imposed to merge these features and pass them through an $MLP$ layer as a regressor and reshape to 2D to generate the final IR drop prediction map.
$z^{prediction}_{1d}\ =\ MLP(y^{gnn}\ \ \|\ \ y^{cnn}),$ (8a)
$z^{prediction}_{2d}\ =\ Reshape(z^{prediction}_{1d}),$ (8b)
@@ III-D Discussion
A previous work [21] on static IR drop prediction also highlighted the importance of PDN structures in achieving IR drop predictions. The authors transformed the PDN structure into a PDN density map (each pixel denotes the total number of PDN strips covering this location) and an effective distance map from the power supply pins. Then, a fully convolutional network, IREDGe, utilizing these additional features for IR drop map generation. In contrast, our paradigm manifests an obvious difference with this work. Firstly, PDNGraph represents the PDN structure using a directed graph. We take into account the specific locations of the PDN on the layout and bridge finely cell-PDN physical relation during the graph construction. In cases where the PDN of a chip is relatively uniform, the PDN density map would become a uniform map which powerless to capture the PDN. Secondly, the edges direction in PDNGraph simulating the current flow from the PDN structure to the cells. The graph branch in PDNNet can favorably perceive and directionally aggregate the features of neighboring features, which cannot be achieved by CNNs.